### v2

# ZeroFlow: Overcoming Catastrophic Forgetting is Easier than You Think

Abstract
Backpropagation provides a generalized configuration for overcoming catastrophic forgetting. Like, SGD and Adam are commonly used for weight updates in continual learning and continual pre-training. In practice, permission to access gradient information is not always granted (the gradient ban), such as black-box APIs, hardware limitations, and non-differentiable systems. To bridge this gap, we introduce the first benchmark ZeroFlow to evaluate gradient-free optimization algorithms for overcoming forgetting. This benchmark examines a suite of forward pass methods across multiple methods, forgetting scenarios, and datasets. We find that forward passes alone are enough to overcome forgetting. Our findings reveal new optimization principles that highlight the potential of forward-pass in mitigating forgetting, managing task conflicts, and reducing memory demands, alongside novel enhancements that further mitigate forgetting with just one forward pass. This work provides essential insights and tools for advancing forward pass methods to overcome forgetting.
1 Introduction
Catastrophic forgetting remains one of the major challenges on the path to artificial general intelligence (AGI) [21, 58, 40, 67], i.e., models tend to forget previously learned tasks when trained on new ones on time-evolving data flow [21, 8, 13]. This phenomenon is commonly seen across various tasks [13], including continual learning (CL) [67, 58], fine-tuning of foundation models (FMs) [39, 71, 63], and continual pre-training (CPT) [37, 49, 72], etc. Among them, optimization algorithms play a crucial role, e.g., SGD has become the default choice during CL [57, 51], while Adam is frequently seen in fine-tuning foundation models [37, 71]. These optimization algorithms in tandem with various methods (ranging from regularization and rehearsal strategies to architectural changes) rely on gradient information to avoid forgetting [68, 3, 36, 24]. Nonetheless, in real-world scenarios, gradient information is not always available or computable (i.e., the gradient ban), like,
Figure 1: Illustrations of ZeroFlow. New tasks (or downstream tasks) arrive sequentially, the gradient bans block the model from learning and memorizing using backpropagation. Our ZeroFlow overcome this issue via forward passes.
Scenario i: large language models as a service (LLMaaS) and black-box APIs.
Scenario ii: hardware systems that do not support principled backpropagation.
Scenario iii: AI for science with non-differentiable underlying systems.
In other words, Scenario i implies that pretrained models are monetized [10, 52, 42] (model owners do not publicly release their pretrained models but instead the service), i.e., only the inputs and outputs are accessible [18, 53]. Scenarios ii/iii implies that the limitations prevent or restrict the execution of backpropagation [32, 59], i.e., extremely high memory demands [39, 24], unsupported systems and hardware [25], or non-differentiable functions, etc [55, 20]. The above means that typical methods for overcoming forgetting are not available because backpropagation is banned, as Figure 1. This yields the primary question to be explored,
(Q) Could we establish a benchmark under gradient bans for overcoming catastrophic forgetting, and explore the overlooked optimization principles?
(a) EASE on average accuracy
(b) EASE on forgetting
(c) APER on average accuracy
(d) APER on forgetting
Figure 2: ZeroFlow Evaluation Results of Catastrophic Forgetting. We visualize the evaluation results of 2 models in several ZeroFlow dimensions (average accuracy over all tasks and a forgetting metric). For comprehensive numerical results, please refer to Table 1.
To tackle (Q), a natural idea is to use the forward pass method [23, 1, 44] instead of backpropagation to overcome forgetting. Such as the zeroth-order (ZO) optimization method [15, 43, 38, 19], category of methods is well-suited to this issue due to their relaxed information requirements (requiring only the function values rather than gradients). Under gradient bans, DECL and DFCL first attempt to overcome forgetting from a stream of APIs [61], but they focus on synthetic data level rather than optimization. Therefore, it remains elusive whether benchmark studies using gradient-free methods can mitigate catastrophic forgetting.
In this work, we explore several Zeroth-order optimization methods on dynamic data Flow (as shown in Figure 1), examining their performance across various forgetting scenarios, model types, and evaluation metrics. Through a detailed analysis, we reveal the overlooked potential of forward passes and various ZO methods in overcoming catastrophic forgetting. This benchmark study offers an easier way to overcome forgetting and helps reveal the pros and cons of these methods in alleviating forgetting. Extended from the gained insights, we introduce a new improvement known as the periodic gradient technique, which further improves ZO optimization for overcoming catastrophic forgetting. Simply put, we can better and more efficiently mitigate forgetting by using just one forward pass.
Our rationale for choosing the ZO optimization algorithms to overcome forgetting for the following two key considerations: (i) implementation cost minimization, that is, we expect minimal modifications to existing optimizers. (ii) theory of diversity, that is, we expect to cover diverse optimization methods. These considerations ensure that our benchmark is comprehensive and simplified. And, an appealing property is that we need only forward passes to be enough to overcome forgetting. Maybe, once is all it takes!
To sum up, our contributions are listed below,
(i) We propose the first benchmark ZeroFlow for overcoming forgetting under gradient bans. This benchmark includes our investigations into 7 forward pass optimization algorithms, several forgetting scenarios and datasets with varying complexity, and task sequences (as Figure 4).
(ii) Through this benchmark, we uncover overlooked optimization principles and insights into how forward pass can mitigate forgetting. These include the role of forward pass in managing task conflicts and the trade-offs between forgetting and memory efficiency. We proved that catastrophic forgetting can be overcome in an easier way!
(iii) Apart from a comprehensive evaluation of catastrophic forgetting, we introduce two enhancement techniques, which further improve the performance and efficiency of just one forward pass to overcome forgetting.
2 Literatures
2.1 Catastrophic forgetting
Catastrophic forgetting occurs across various tasks, including CL, fine-tuning of FMs, and CPT [67, 58, 71, 37]. To mitigate this issue, various methods have been proposed [45, 26, 54, 30]. In CL, methods range from regularization and rehearsal strategies to architectural changes [28, 3, 36]. Lately, pre-trained models (PTM) further advanced these methods due to their strong generalization [62, 12], as seen in PTM-based CL [69]. All these methods share a common goal: achieving an optimal balance between learning plasticity and memory stability [58]. In FMs, catastrophic forgetting often arises from overfitting to small fine-tuning datasets during CPT or fine-tuning [37, 71]. Common techniques to address this include learning rate adjustment, parameter-efficient fine-tuning, mixed data strategies, and instruction tuning [37, 24]. Additionally, as foundational models increasingly gain multimodal capabilities, the complexity of catastrophic forgetting also intensifies [71].
2.2 Optimization for catastrophic forgetting
Two broad categories of optimization methods exist for overcoming forgetting, (i) Standard Optimization. SGD and the Adam family are frequently employed to investigate catastrophic forgetting [4, 21, 40]. For instance, in CL, various CL methods predominantly utilize the SGD optimizer for standard evaluations [57, 51, 70]. In fine-tuning the LLM, the Adam series is commonly used to observe forgetting phenomena [37, 71]. Some works explored orthogonal spaces with these standard optimizers to alleviate forgetting [35, 14, 47], such as OGD [11], and GPM [47]. Moreover, other works [11, 5, 35] modified the gradients in the standard optimization process to align the learning spaces of new and old tasks, such as Uni-Grad [30]. The core of these efforts [9, 48] is to find an equilibrium between learning and forgetting in optimization. (ii) Sharpness-aware Optimization. This series of methods [22, 16, 65, 73] started to draw attention due to the effectiveness of the flat minimum in mitigating forgetting [30, 29, 4, 41]. For instance, FS-DPGM [9], F2M [48], DFGP [60], SAM-CL [56], etc. overcome forgetting in the flatness areas of different configurations. C-Flat [3] proposed a CL-friendly general optimization framework, that holds promise as a baseline optimizer for overcoming catastrophic forgetting.
Our work. The works mentioned above are all rooted in a gradient feedback mechanism. Such mechanisms are powerless against catastrophic forgetting without explicit gradient information. Our work overcomes forgetting only via forward pass instead of gradient feedback.
3 Exploring Zeroth-order Optimization to Overcome Forgetting
3.1 Zeroth-order Optimization
Zeroth-order (ZO) optimization has been extensively studied over the years within the realms of numerical computation and approximation algorithms. It functions as an alternative solution for estimating descent directions in scenarios where first-order (FO) gradients are either inaccessible or infeasible to compute. Considering a deep learning model parameterized with $\mathbf{\theta}\in\Theta\subseteq\mathbb{R}^{d}$, and given a mini-batch $\mathcal{B}$ extracted from the training dataset $D=\{(x_{i},y_{i})\}_{i=1}^{m}$. Let $L(\mathbf{\theta};\mathcal{B})$ denote the empirical loss, then the genetic formulation of ZO optimization follows Algorithm 1.
1 Input Initialized model parameters $\mathbf{\theta}_{0}\in\Theta\subseteq\mathbb{R}^{d}$, training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}\in\mathcal{X}\times\mathcal{Y}$, empirical loss function $\mathcal{L}$, learning rate $\eta_{t}$, gradient perturbation vector $\mathbf{\xi}$, and descent direction computation $\phi(\cdot)$
2 while $\theta_{t}$ not converged do
3       Sample mini-batch $\mathcal{B}$ from $\mathcal{D}$
4       Step 1. $\triangleright$ ZO gradient estimation:
5       $\hat{\mathbf{g}}_{t}=\hat{\nabla}\mathcal{L}(\theta,\mathbf{\xi};\mathcal{B})$
6       Step 2. $\triangleright$ Descent direction computation:
7       $\mathbf{h}_{t}=\phi\left(\{\hat{\mathbf{g}}_{i}\}_{i=1}^{t}\right)$
8       Step 3. $\triangleright$ Parameter updating:
9       $\mathbf{\theta}_{t+1}=\mathbf{\theta}_{t}-\eta_{t}\cdot\mathbf{h}_{t}$
10       $t=t+1$
Output Updated model $\mathbf{\theta}_{t}$
Algorithm 1 Genetic formulation of ZO optimization
1) ZO gradient estimation. Randomized Gradient Estimation (RGE [43]) and Coordinate-wise Gradient Estimation (CGE [2]) perturb the model using $\mathcal{\xi}$, which is generated either from a random unknown distribution (in RGE) or by modifying individual coordinates (in CGE), and then observe the changes in the loss function $\mathcal{L}$ after each perturbation, step by step, to provide a reliable gradient estimate. However, due to their reliance on slow single-direction perturbation, these methods are not well-suited for deep learning tasks, as performing a full perturbation in high-dimensional parameter spaces is time-consuming. For instance, the ImageNet dataset has input dimensions of 224×224×3, totaling approximately 0.13 million parameters, and rendering over each dimension iteratively leads to query inefficiency. The standard Simultaneous Perturbation Stochastic Approximation (SPSA) improves efficiency by generating pairs of symmetric vectors and perturbing in multiple directions simultaneously, as follows
^$\hat{\nabla}L(\theta,\xi;\mathcal{B})=\frac{f(\theta+\epsilon\xi;\mathcal{B})-f(\theta-\epsilon\xi;\mathcal{B})}{2\epsilon}\xi^{-1}.$ (1)
Where $\epsilon$ is a positive scaler and $\xi$ is recommended to follow a symmetric distribution with finite inverse moments (e.g., the Rademacher distribution). The symmetric distribution ensures unbiased exploration of perturbations in both positive and negative directions of parameters at each step. And the finite inverse moments property guarantees that the steps are well-controlled, avoiding excessively large steps due to $\xi^{-1}$ drawn from the distribution (e.g., $\mathbb{E}[1/|\xi|^{p}]$ for some large p), which would otherwise lead to an unstable optimization process. In practical implementations for models with a large number of parameters (e.g., MeZO [38] in large language models (LLMs)), Gaussian noise with zero mean induces substantial perturbations, thereby enhancing exploration across the parameter space and facilitating the escape from local minima. This methodology achieves gradient estimation with only two objective function evaluations, rendering its computational cost independent of input dimensionality. Such computational efficiency has established SPSA as a preferred method for addressing the complexities of high-dimensional deep learning tasks. While increasing $q$ in $q$-SPSA can improve stability in the update direction, setting $q=1$ is sufficient for pretrained LLMs  [38].
2) Descent direction computation. In unconstrained optimization for deep learning, the last gradients $h_{t}$ generally coincide with the estimated ZO gradients $\hat{g_{t}}$ (e.g., ZO-SGD [19], ZO-SCD [31]). To reduce approximation errors, ZO-SGD-sign [34] applies an element-wise $sign(\cdot)$ operation. Additionally, ZO-SVRG [33], inspired by variance reduction methods in first-order optimization, adjusts the update step by using estimated gradients from previous training examples. The CARS method [27] adaptively selects the smallest function value in each iteration, which helps maintain monotonicity in the optimization process.
3) Parameter updating. Normally, for most ZO methods, parameters are updated in a similar way with FO optimizers, and the learning rate $\eta_{t}$ is set to constant. Except for the special design for achieving some constraint prerequisites, several methods make an effort to strike a balance between converge speed and accuracy. ZO-AdaMM [7] uses an adaptive learning rate and refines gradient estimation by incorporating momentum from past information. This approach is particularly effective in handling complex and evolving optimization landscapes, where the function’s behavior may vary over time or be hard to capture with straightforward gradient approximations.
3.2 Zeroth-order Optimization for Catastrophic Forgetting
Rationality. ZO optimization leverages the function values of the forward passes to approximate FO gradients, making it feasible to avoid gradient bans. This feature enables seamless integration into common forgetting scenarios, such as continual learning. We explore it in the following three categories.
i) Memory-based methods maintain a repository of exemplars from previous tasks and dynamically adjust the overall loss function by combining these stored samples with new data based on learning progress.
$\mathcal{L}_{total}=\frac{1}{N_{context}}\mathcal{L}_{cur}+(1-\frac{1}{N_{context}})\mathcal{L}_{replay},$ (2)
where $N_{context}$ represents the number of contexts encountered so far. In Experience Replay [46], both components use classification loss based on their respective data distributions, so ZO gradients can be expressed as ^$\hat{\nabla}\mathcal{L}_{cur}$ and ^$\hat{\nabla}\mathcal{L}_{replay}$ respectively. However, in the emerging generative replay workflows [50], Eq. 2 may introduce additional loss for the training of generators. In this case, the generator can be trained using standard backpropagation or in conjunction with ZO training without FO gradients.
ii) Extension-based methods can be divided into fixed and dynamic architectures. Fixed architectures separate model parameters for specialized context learning, while dynamic architectures expand the model size during adaptation. Both approaches mitigate catastrophic forgetting from the model’s perspective and enable model-agnostic ZO solutions.
iii) Regularization-based methods penalize significant changes to parameters important for old tasks or maintain the output distribution with respect to previous inputs. The template loss function is given by
$\mathcal{L}_{total}=\mathcal{L}_{cur}+\alpha\mathcal{L}_{reg},$ (3)
where $\alpha$ is a coefficient hyperparameter. The FO gradients from dual objectives ($\mathcal{L}{cur}$ for adaptation and $\mathcal{L}{reg}$ for preservation) drive optimization toward their respective optima, achieving inter-task equilibrium. Notably, ZO gradient estimates, though obtained in a noisy environment, exhibit comparable optimization behavior.
(a) FO-SGD
(b) ZO-SGD
Figure 3: The Trajectory of FO and ZO Optimization during Overcoming Forgetting. The trajectory has taken when using the total loss from both tasks (cyan) and the gradients from each individual task at fixed points during optimization (red and blue). The trends of ZO optimization hold the potential to manage forgetting and learning.
As shown in Figure 3, we visualize and compare the optimization trajectories of ZO and FO methods in the learning-memory trade-off dynamics of continual learning. The remarkable similarity between these trajectories motivates our investigation, suggesting ZO optimization’s potential in effectively managing the balance between learning and forgetting.
Potential. The intrinsic optimization mechanism of ZO exhibits particular promise in continual learning scenarios. Intuitively, ZO perturbs parameters using random or coordinate-wise directional vectors and observes changes in the evaluation function, effectively optimizing within a noisy environment. This approach enables small parameter modifications to yield significant impacts on target objectives, resulting in distinctive gradient estimations compared to FO optimization. Notably, while ZO methods do not explicitly incorporate sharpness regularization terms [17], they naturally facilitate the exploration of flat regions in parameter space. The influence of optimizing flat regions with ZO approaches in continual learning can be summarized in two main manifolds: (i) For previous tasks, the noise-induced parameter robustness enhances resilience against perturbations from new task adaptation; (ii) For new tasks, empirical evidence suggests that convergence to flat minima generally leads to lower generalization error.
Table 1: ZeroFlow Evaluation Results on CIFAR-100, ImageNet-A, CUB and OmniBenchmark. This table compares the average accuracy, final accuracy, and forgetting measures of two models, and four forgetting scenarios. More intuitive trend please see Figure 4. Note that all ZO optimization w/ the query budget $q=1$, which means that all our evaluations use just one forward pass.
Method Optimizer Strategy CIFAR-100 CUB ImageNet-A OmniBenchmark
Avg Last Fgt Avg Last Fgt Avg Last Fgt Avg Last Fgt
EASE SGD FO 91.23 85.96 7.32 89.31 83.76 9.61 61.24 51.02 10.84 74.73 67.40 15.11
ZO 78.62 68.40 15.64 88.94 82.91 8.08 57.87 48.32 11.08 73.50 66.60 17.78
Sign 83.21 75.88 10.58 89.81 84.61 8.10 59.15 49.31 11.77 73.81 66.75 17.21
Conserve 82.22 75.88 8.93 89.21 83.42 10.31 58.61 48.58 12.41 77.07 70.73 14.87
Adam FO 90.56 84.82 7.69 84.44 77.10 10.51 59.60 47.20 19.08 74.27 66.28 15.63
ZO 83.36 76.09 10.16 89.49 84.14 8.67 58.90 48.72 12.35 76.15 69.69 15.87
Sign 83.14 76.01 10.44 89.82 84.65 8.21 58.97 48.85 12.20 77.12 71.08 14.68
Conserve 82.15 75.65 9.24 89.82 84.61 8.40 59.23 48.85 12.81 77.19 70.99 14.68
- Forward 82.26 76.05 8.74 89.26 83.67 9.35 57.76 48.19 11.03 77.00 70.74 14.99
APER SGD FO 82.31 76.21 7.33 90.56 85.16 5.19 59.50 49.37 9.91 78.61 72.21 7.87
ZO 82.33 76.21 7.36 90.53 85.20 5.12 59.58 49.51 10.02 78.60 72.21 7.85
Sign 82.32 76.23 7.32 90.42 85.28 4.96 59.65 49.77 9.89 78.60 72.26 7.78
Conserve 82.31 76.21 7.33 90.62 85.28 5.05 59.68 49.70 10.18 78.61 72.21 7.87
Adam FO 82.31 76.21 7.33 90.56 85.16 5.19 59.60 49.77 10.06 76.60 72.21 7.85
ZO 82.12 75.45 7.47 90.33 84.31 6.01 58.89 49.24 9.32 78.44 72.10 7.87
Sign 82.01 75.60 7.38 89.86 84.18 5.99 57.82 48.12 9.72 78.26 72.05 7.75
Conserve 82.21 75.98 7.34 89.96 84.48 5.90 57.86 47.53 10.00 78.61 72.21 7.87
- Forward 82.32 76.22 7.32 89.47 83.38 6.24 58.25 47.99 9.62 77.61 71.45 7.87
Risk. Although ZO demonstrates superior generalization abilities, its practical performance is limited by optimization strategies and the complexity of the optimization setting. Despite significant efforts to reduce convergence error, optimizing models from scratch in high-dimensional space remains challenging due to slow convergence speed (proportional to the parameter dimension $d$). For instance, origin CGE-based ZO training for a model with 12k parameters takes 70.32 hours in DeepZero [6]. Such computational demands render from-scratch training impractical for high-dimensional continual learning models, particularly those employing expansion-based architectures. Consequently, we focus our discussion on leveraging ZO optimization to overcome forgetting within a pre-training context.
4 ZeroFlow Benchmark
This section delves into the empirical performance of ZO optimization in overcoming catastrophic forgetting. Our ZeroFlow includes evaluating average performance along incremental stages, last stage accuracy, forgetting, and efficiency, considering the complexity of different datasets, and considering various model types.
4.1 Benchmark Setups
Forgetting scenarios, schemes, and models. We assigned the evaluation to a common catastrophic forgetting scenario, which is continual learning. We focus on four fine-tuning tasks, considering their complexity from the typical dataset CIFAR and CUB to dataset ImageNet-A, and OmniBenchmark, which has a large domain gap. For the method used on these tasks, we explore two SOTA schemes: EASE and APER. Both models are initially configured with ViT-B/16-IN21K and subsequently fine-tuned on four datasets. Following [66], each dataset is split equally into 10 tasks by class. For example, OmniBenchmark contains 300 classes, with 30 classes learned in each incremental stage. No memory is allowed for storing past examples.
Benchmark setup and details. To evaluate the application of ZeroFlow in forgetting scenarios, we include the methods described in Sec. 3.1, specifically ZO [19, 38], Sign [34], and Conserve [27, 64], in comparison with their FO counterparts using SGD and Adam optimizers [7]. Additionally, as highlighted in [64], Forward-Grad [44] which relies on forward mode automatic differentiation, potentially becomes a missing but competitive forward pass baseline. In a nutshell, ZeroFlow covers 7 forward pass methods: ZO-SGD, ZO-SGD-Sign, ZO-SGD-Conserve, ZO-Adam, ZO-Adam-Sign, ZO-Adam-Conserve, Forward-Grad.
Evaluation metrics. Overall, we use two sets of metrics as ZeroFlow evaluation measures: accuracy and efficiency. Accuracy measures include average accuracy over all tasks, the last-task accuracy, and a forgetting metric. Efficiency measures include memory efficiency (GPU usage), query efficiency, and runtime efficiency. Together, these metrics provide insights into the resources required for ZO optimizations to overcome forgetting.
Table 2: The Memory Cost (in GB) and the Runtime Cost (in seconds) of Each Optimizer on Four Forgetting Scenarios. The per-epoch runtime in seconds (s). ZO-SGD w/ the query budget $q=1,4$ and all other optimizers w/ the query budget $q=1$.
Optimizer Memory $\Downarrow$ CIFAR-100 CUB ImageNet-A OmniBenchmark
FO-SGD 12.08 GB 59.3s 16.1s 12.2s 109.6s
ZO-SGD ($q=1$) 2.41 GB 32.4s 8.3s 6.8s 66.8s
ZO-SGD ($q=4$) 2.41 GB 111.7s 28.7s 18.0s 221.3s
ZO-SGD-Sign 2.41 GB 32.4s 8.3s 6.8s 66.8s
ZO-SGD-Conserve 2.41 GB 70.1s 15.7s 12.4s 133.7s
Forward-Grad 3.94 GB 45.9s 11.1s 9.0s 82.0s
(a) EASE on last accuracy
(b) APER on last accuracy
Figure 4: ZeroFlow Evaluation Results for Forgetting. We visualize the evaluation results of 2 models in the last-task accuracy.
4.2 Evaluation Results of ZeroFlow
ZeroFlow evaluation on continual learning. In Table 1, we evaluate the performance of different BP-free and BP-based (FO-SGD and FO-Adam) methods in a typical forgetting scenario (continual learning). We use two SOTA models as examples and investigate SGD and Adam optimizers, seven forward pass methods, and four commonly used datasets. Several observations are listed below,
First, the performance of the ZO method is comparable to or even surpasses that of the FO method across almost all forgetting metrics and datasets. However, as will be shown later, the FO method requires significantly more memory overhead. This suggests that forward passes alone can effectively mitigate forgetting, and the ZO method offers a simpler, more efficient alternative. In some cases, such as with ZO-Adam and ZO-SGD on OmniBenchmark, ZO methods even outperform FO methods.
Second, the Forward Grad method demonstrates competitive performance when compared to other ZO and FO methods. Unlike typical ZO methods, Forward Grad utilizes a unique forward pass mechanism, making it a promising baseline for future studies. A more intuitive trend in overcoming forgetting can be observed in Figure 4. These observations motivate further exploration into the effectiveness of the ZO method.
ZeroFlow helps manage memory and runtime. In Table 2, we compare the efficiency of various ZO and FO optimizers in overcoming catastrophic forgetting, focusing on two main aspects: memory cost (in GB) and runtime cost (in seconds). First, naïve ZO optimization reduces memory cost by approximately fivefold compared to FO optimization. Meanwhile, ZO optimization decreases runtime cost per iteration by around 50% compared to FO, which significantly broadens the applicability of ZO optimization in overcoming forgetting. Second, ZO and sign methods show similar levels of efficiency in terms of memory efficiency and runtime costs. Although the number of queries can impact the efficiency of ZO methods, it does not destroy memory management efficiency. Third, Conserve also features efficient memory management, though its runtime cost is about 2x longer than the ZO. This may explain why Conserve performs better in a few cases, as shown in Table 1.
Figure 5: Performance Comparison under Different Query Mumbers. Both optimizers show improved performance as query numbers increase.
(a) FO-Adam
(b) ZO-Adam
(c) ZO-Adam ($q=4$)
(d) ZO-Adam-Sign
(e) ZO-Adam-Conserve
(f) FO-SGD
(g) ZO-SGD
(h) ZO-SGD ($q=4$)
(i) ZO-SGD-Sign
(j) ZO-SGD-Conserve
Figure 6: The Trajectory of Different Optimization during Overcoming Forgetting. Red denotes the minimum of new task, blue denotes the minimum of old task. The trajectory taken when using the total loss from both tasks (cyan).
Trade-off between performance and query number. As shown in Figure 5, we investigate the impact of query numbers on optimization performance, comparing SGD and Adam optimizers in the zeroth-order setting. Both optimizers demonstrate improved performance as query numbers increase across [1,2,4,8,16,32], suggesting that additional function evaluations enable more accurate gradient estimation. The results suggest that in scenarios where function evaluation costs are manageable, higher query numbers can yield substantially better performance, with Adam being particularly effective at leveraging the additional gradient information for enhanced optimization outcomes.
5 Insights and Discussions
As shown in Figure 6, we visualized the optimization trajectories of both forward pass and backpropagation methods. Our analysis reveals several key insights:
Convergence behavior across optimizer families. In Figure 6, both FO and ZO methods demonstrate successful convergence to the minima of new and old knowledge spaces, regardless of whether they use Adam or SGD as their base optimizer. This convergence consistency validates our theoretical foundation.
Distinct trajectory characteristics of FO and ZO. FO approaches (Figure 6, 6) show smoother optimization paths due to their access to exact gradient information. In contrast, ZO methods demonstrate varying degrees of exploration behavior through trajectory jitter. This exploration pattern is particularly pronounced in ZO-Adam variants compared to ZO-SGD variants, indicating that the base optimizer choice significantly influences the exploration-exploitation trade-off during optimization.
Path characteristics in ZO optimization. Comparing base ZO methods with their $q=4$ counterparts (Figure 6 vs 6, Figure 6 vs 6), we observe that increasing query numbers leads to smoother trajectories, suggesting that more queries help provide more stable gradient estimates. The Sign variants (Figure 6, 6) demonstrate more pronounced oscillations in their trajectories, particularly visible in the ZO-Adam-Sign case. In contrast, the Conservative variants (Figure 6, 6) maintain relatively stable paths that better balance between the old and new task minima.
Distinct characteristics between optimizer families. Adam-based approaches (Figure 66) demonstrate more oscillatory trajectories with frequent direction adjustments, indicating a more dynamic exploration of the loss landscape. In contrast, SGD-based methods (Figure 66) exhibit smoother and more stable trajectories, suggesting a more gradual progression toward the optimization objective. These distinct optimization patterns could influence how each method balances between preserving old task knowledge and adapting to new tasks.
6 New Enhancement to Mitigate Forgetting
In this section, we further investigate several techniques to stabilize the updates for mitigating forgetting. In ZO optimization, the estimation of the gradients relies on a finite difference of the objective function (i.e., the gradients are estimated by querying the function values). We set $q=1$ in $q$-SPSA for efficiency in the previous benchmark settings. However, limited queries cannot capture the accurate ZO directions, leading to large variances in the gradient estimation. In continual learning scenarios where models learn tasks sequentially, the high variance inherent in ZO gradient estimation poses a critical challenge. This uncertainty, coupled with large parameter updates, can significantly amplify catastrophic forgetting through undesirable perturbations to previously learned knowledge. This issue is amplified for larger models such as ViTs and model-expansion architectures, due to their high-dimensional parameter spaces. Though increasing query numbers can stabilize the gradient estimates, it leads to prohibitive computational costs in continual learning settings. Thus, exploring variance-reduced optimization algorithms is crucial for practical ZO-based continual learning. Specifically, we propose two complementary approaches to stabilize the ZO optimization process while maintaining computational efficiency:
(a) FO-SGD
(b) ZO-SGD
Figure 7: The Variation in Function Values of the Forward Passes. Function values for new tasks is highlighted in red, old tasks is highlighted in green.
First, leveraging historical gradient information to overcome forgetting. When learning new tasks, models leverage previously learned parameters while prioritizing the preservation of crucial parameters for old tasks. To mitigate interference from new tasks, we propose reweighting old task gradients with historical gradients, which can stabilize perturbations caused by low query loops in ZO optimization. Figure 7 illustrates the function value trajectories for both old and new tasks. While FO optimization shows smooth convergence toward the global optimum, ZO optimization exhibits a more volatile path. Notably, objectives related to old tasks demonstrate smaller changes in both magnitude and variance. This observation motivates us to stabilize the optimization by reducing changes to old gradients through a linear combination with historical gradients: $g_{old}=(1-\alpha)g_{old}+\alpha g_{historical}$, where larger $\alpha$ indicates greater reliance on historical information for stability, at the cost of reduced contrast with new task gradients.
Table 3 validates the effectiveness of historical estimation in mitigating catastrophic forgetting. Modest proportions of historical information (e.g., 20%, 40%, 60%)) outperform ZO-SGD (0%), effectively controlling perturbations while maintaining a low query budget ($q=1$). While this approach helps stabilize updates, the inherent noise in ZO estimation requires additional consideration, leading us to explore complementary variance reduction techniques.
Table 3: Effectiveness of Historical Estimation in Mitigating Forgetting. The proportion of 0% denotes that the plain optimizer ZO-SGD.
Metrics Proportion
0% 20% 40% 60% 80%
Avg 57.87 58.90 58.76 58.34 57.83
Last 48.32 49.04 48.84 48.42 48.10
Fgt 11.08 11.79 11.78 11.60 11.57
Table 4: Effectiveness of Sparsity-induced Estimation in Overcoming Forgetting. The proportion of 0% denotes that the plain optimizer ZO-SGD.
Metrics 0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Avg 57.87 59.17 59.46 59.29 59.39 59.45 59.26 59.39 59.38 59.47
Last 48.32 48.58 49.05 48.72 48.91 49.24 49.11 49.05 49.11 49.24
Fgt 11.08 12.65 12.17 12.76 12.53 12.37 12.36 12.54 12.46 12.33
Second, sparsity-induced estimation helps to overcome forgetting. Gradient sparsification reduces the computational cost and memory usage by setting parts of the gradients to zero or by retaining a subset of the gradient components. In ZO optimization, the gradients for new tasks are often highly uncertain due to the approximation nature of the gradient estimation. To reduce this variance, we implement random sparsification by creating a seed-based mask and setting gradients outside the mask to zero. By reducing the number of non-zero gradient components, we aim to stabilize the optimization process and mitigate the noise in gradient updates. However, it is important to balance sparsity carefully, as excessively high sparsity could impair the learning of new tasks and exacerbate catastrophic forgetting, especially when too many important gradient components are zeroed out. The sparsity level is varied in our experiments, ranging from 10% to 90%.
Table 4 shows the performance of sparsity-based ZO optimization in overcoming catastrophic forgetting. We observe that the sparse technique improves the average and last accuracy across all scales, which implies that forgetting is effectively controlled. The reduction in volatility can be attributed to the sparse strategy yielding smoother gradient estimates compared to plain ZO-SGD, effectively bounding variance to a low level and thus mitigating forgetting. Moreover, the robust performance across different sparsity ratios provides strong evidence for the efficacy of variance control in addressing forgetting.
7 Conclusion
This paper introduces ZeroFlow, a benchmark study that probes a series of forward pass methods for overcoming catastrophic forgetting. This work resorts to an easier way (no need for backpropagation and activation storage) to overcome forgetting. Concretely, our benchmarks include various forward pass methods, forgetting scenarios, and evaluation metrics. We also reveal the overlooked optimization principles for overcoming forgetting via forward pass. Based on these insights, we propose two easier and better enhancement to overcome forgetting and extend the application of related methods easily.